OpenStudy (anonymous):

1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4.

OpenStudy (kymber):

Plug the values into y = mx + b and solve for b

OpenStudy (anonymous):

can you show me an example

OpenStudy (kymber):

Parallel lines have same slope, So you just put 4 in as m & 5, -2 in as x & y then you solve it.

OpenStudy (anonymous):

so it would be -2 = 4*5b

OpenStudy (kymber):

Yupp! then you solve it, & plug m & b back into the y= mx + b equation and there's your answer.

OpenStudy (anonymous):

i am still confused

OpenStudy (kymber):

Solve this. What did you get?